# FaceQSORT: a Multi-Face Tracking Method based on Biometric and Appearance Features

### v2

@@ Abstract
Tracking multiple faces is a difficult problem, as there may be partially occluded or lateral faces. In multiple face tracking, association is typically based on (biometric) face features. However, the models used to extract these face features usually require frontal face images, which can limit the tracking performance. In this work, a multi-face tracking method inspired by StrongSort, FaceQSORT, is proposed. To mitigate the problem of partially occluded or lateral faces, biometric face features are combined with visual appearance features (i.e., generated by a generic object classifier), with both features are extracted from the same face patch. A comprehensive experimental evaluation is performed, including a comparison of different face descriptors, an evaluation of different parameter settings, and the application of a different similarity metric. All experiments are conducted with a new multi-face tracking dataset and a subset of the ChokePoint dataset. The ‘Paris Lodron University Salzburg Faces in a Queue’ dataset consists of a total of seven fully annotated sequences (12730 frames) and is made publicly available as part of this work. Together with this dataset, annotations of 6 sequences from the ChokePoint dataset are also provided.
@@ 1 Introduction
Multiple object tracking (MOT) [1] is about finding the trajectories of all objects appearing in a sequence. These objects could be from different classes or a single class (e.g. pedestrians, cars, persons, etc.). When considering faces, MOT is referred to as multi-face tracking. This work focuses on tracking the faces of people moving towards a gate in a queue.
@@
The assumed scenario is that a group of people is moving towards a gate (e.g., to enter a sports stadium). As they move towards a single gate, they form a queue. However, as this is a leisure activity, this queue will not be very well organized (people chat with each other, they eat, they move around in a disorderly fashion, there may be pushing and shoving, etc.). In the assumed scenario, the main objective is to track people’s faces as they move towards the gate. For this purpose, a camera is mounted on the gate. The camera is directed at the queue such that the first person visible is standing directly in front of the gate. However, since the queue is probably not very well organized, the main tracking challenges are occlusions (partially and full), out-of-plane rotations (lateral faces) and non-linear motion.
To solve a general MOT problem, basically two paradigms exist: (i) joint-detection-association (JDA), and (ii) tracking-by-detection (TbD). With a JDA based tracker, basically everything is learned (end-to-end), i.e., the object detection, the relevant cues for the association and the association of objects across frames. Recent methods in this context are [2, 3, 4, 5, 6]. In this work, however, the focus is on TbD methods [7, 8, 9, 10, 11, 12]. The first step of a tracker based on the TbD paradigm is the detection of all objects (e.g., faces) in the current frame. The tracking performance is therefore directly related to the detection performance. In [10], a Spatial-Temporal Topology-based Detector (STTD) is proposed. A feature mixing strategy is proposed in [13] to improve the detection performance in the context of MOT. Tracking is then about finding bipartite matches for all detected objects and active tracks (i.e., association problem). A common way to solve this association problem is to apply the Hungarian algorithm [14] to a cost matrix (representing the costs of matching the $i$-th active track with the $j$-th detected object). These costs can, for example, be based on motion cues and/or appearance cues. Prominent examples for TbD tracker are Simple Online Realtime Tracking (SORT) [15] (motion based association), DeepSORT [16] (associated based on motion and appearance features) and StrongSORT [17] (improved version of DeepSORT). In [12], a transformer is used to estimate complex motion cues. A combination of motion and appearance models in a single network, called UMA, is proposed in [18]. In [9], a fusion of six distance metrics (for motion and appearance cues) is used for association. One advantage of TbD trackers over JDA trackers is that pre-trained models can be utilized for the individual subtasks (e.g., a face detector and face recognition model), while performance is competitive (as demonstrated in [19]). A form of TbD trackers are multi hypothesis trackers [20, 21]. In multi hypothesis tracking, several track hypothesis are maintained simultaneously for the trajectory of each object.
In TbD based multi-face trackers, association is based on (biometric) face features (i.e., deep neural networks trained in the context of face recognition). For example, a two-stage association based on motion and face features (ArcFace [22] trained on the MS-Celeb-1M dataset [23]) is proposed in [24]. Multi-face tracking based on DeepSORT [16] is proposed in [25]. The authors evaluate different combinations of face detectors (i.e. MTCNN [26], SSD [27] and R-FCN [28]) and a face model trained with two different loss functions (i.e., cosine softmax classifier loss [29] and angle softmax classifier loss [30]). Face features of the upper facial regions (i.e. eyes, eyebrows and forehead) are used to track masked faces in [31]. A multi-face tracking method based on SORT [15] (i.e., motion based association) with a similarity matching block (comparing the stored with the currently computed faces features (i.e., ArcFace [22])) as a fallback for all unmatched detections showing a frontal face is proposed in [32]. The method is named ReSORT, because IDs can be recovered based on the similarity matching block. Face features as fallback if motion based association fails is also proposed in [33]. In [34], a single object tracking methods to predict the positions of active tracks (motion cues) is used. The position-based matches are then corrected using face features. A multi-face tracking method based on face features only is presented in [35]. Cumulative learning of face features is performed based on a memory module where selectively redundant features are removed. The proposed method is referred as IdOL (Identity Online Learning). A Multi-Camera Face Detection and Recognition (MCFDR) tracking is proposed in [36]. MCFDR combines YOLO [37] face detection with DeepSORT [16], where SphereFace [38] face features are used for association.
The face features used for tracking are usually based on pre-trained face recognition models (e.g., [22, 39, 40, 41]) that are typically trained on more or less frontal, unoccluded faces. In a multi-face tracking scenario, however, lateral faces (out-of-plane rotations) and partially occluded faces occur. To address this issue, a multi-modality tracker, Online Multi-face Tracking with Multi-modality Cascaded Matching (OMTMCM), is proposed in [42]. The proposed method utilizes body and face features and includes two stages: (i) detection alignment, and (ii) detection association. The detected faces and bodies are aligned in the first stage (i.e., detection alignment). Detection association is performed in a cascade, whereby matching based on face features is performed only with those detections that could not be matched with body features. The face features are extracted by a SeNet [43] pre-trained on the VGGFace2 database [44] and the body features by a ResNet [45] pre-trained on the MovieNet database [46]. However, a corresponding body is not always visible for every detected face, especially in the assumed scenario in which people are moving towards a gate in a queue.
@@
In this work, we propose FaceQSORT, a new TbD based multi-face tracker inspired by StrongSORT. Instead of utilizing features extracted from two modalities (e.g., face and body)[42], two different types of features are extracted from the same modality (i.e., detected face). By using two different features from the same image patch, the required presence of body and face, which hardly exists in the considered scenario, is eliminated. The two features (i.e., face features and appearance features) are combined into a single cost value, which forms the basis for association. To the best of our knowledge, FaceQSORT is the first multi-face tracking method that combines two features extracted from the same image patch. All other described multi-face trackers [24, 25, 31, 32, 33, 34, 35] rely solely on face features. The combination of the two features helps to better handle lateral or partially occluded faces. FaceQSORT is evaluated based on a new multi-face tracking dataset. This dataset reflects the assumed scenario and is released as part of this work. To the best of our knowledge, the published dataset is the first multi-face tracking dataset in which people move in an (unorganized) queue towards a camera (gate). There is only one similar dataset, the ChokePoint dataset [47], in which several people move through a portal. In the released dataset, however, a queue is simulated in which people are talking, eating, pushing, shoving, etc.
@@
  • A novel multi-face tracking method (FaceQSORT) is proposed that combines two different features (i.e., biometric (face) features and appearance features) extracted from the same image patch.
  • A comprehensive experimental evaluation is performed comparing different face feature descriptors, analyzing parameter selection and applying a different similarity metric (i.e., Euclidean distance).
  • An ablation study is performed.
@@
The remainder of this paper is organized as follows: the proposed FaceQSORT is described in section 2. Section 3 provides an overview of the proposed dataset. The experimental evaluation is described in section 4. An ablation study is conducted in section 5 and the key insights are summarized in the last section 6.
@@ 2 FaceQSORT
FaceQSORT is a multi-face tracker based on the basic variant of StrongSORT [17]. Since common face recognition models are usually trained on constrained face images, successful association (of detected faces and active traces) of lateral or partially occluded faces based on extracted face features can be difficult. To mitigate this problem, FaceQSORT combines two different types of features extracted from the same image patch (i.e., detected faces). These features are: (i) face features (representing biometric information), and (ii) appearance features (representing visual appearance). The combination of these features is performed on the cost matrix level.
In common TbD based methods, the association relies on a cost matrix $C^{n\times m}$, where $n$ is the number of active trajectories (tracks) and $m$ is the number of detected objects. An element $C_{i,j}$ represents the cost of matching the $i$-th active track to the $j$-th detected object. The main objective of association is to find bipartite matches where the global cost is minimum (i.e., usually solved by the Hungarian algorithm [14]). In principle, $C_{i,j}$ is based on a distance $d$ from a stored/predicted cue of the $i$-th active track to a calculated/observed cue of the $j$-th detected object and is calculated in FaceQSORT as follows,
@@
[…]
The parameter $\lambda\in\mathbb{R}_{[0,1]}$ weights the contribution of the biometric (face) features ($C_{i,j}^{bio}$) and the appearance features ($C_{i,j}^{app}$). Thus, $C_{i,j}^{app/bio}$ represents the similarity in terms of biometric cues and visual appearance from a stored face (active track) and a detected face. A spatial distance between a predicted and an observed position is represented by $C_{i,j}^{pos}$ and weighted by the parameter $\beta\in\mathbb{R}_{[0,1]}$.
The set of detected faces at frame $t$ is denoted by $\Psi$, and $\Psi_{j}^{bio}$, $\Psi_{j}^{app}$ and $\Psi_{j}^{pos}$ represent the extracted biometric features, the extracted appearance features and the observed position (bounding box), respectively. $\Phi$ represents the active track memory, storing biometric features $\Phi^{bio}$, appearance features $\Phi^{app}$ and positions $\Phi^{pos}$ for each active track. A track becomes inactive (is deleted from $\Phi$) if it could not be matched with a detected face in $N_{max}$ consecutive frames. When a new track is added to $\Phi$, it is in a tentative state and is only confirmed if the track could be matched in the next $N_{init}$ frames. Otherwise, the tentative track is deleted from $\Phi$. Similar to StrongSORT, an Exponential Moving Average (EMA) [48] approach is used to update the stored biometric and appearance features from the $i$-th matched track, i.e.,
$\Phi_{i}^{t}=\alpha\Phi_{i}^{t-1}+(1-\alpha)\Psi_{i}^{t},$ (2)
where $\Psi_{i}^{t}$ represents the extracted features from the current frame $t$ of the face associated with track $i$ and the parameter $\alpha$ is a momentum term. As stated in [17], this updating strategy leverages the information of inter-frame feature changes and can mitigate detection noises. $\Phi$ is maintained separately for biometric and appearance features (i.e., $\Phi^{bio}$ and $\Phi^{app}$). The prediction of the position of the $i$-th active track ($\Phi_{i}^{pos}$) in the next frame is based on the NSA Kalman filter [49]. The distance measures used in equation (1) are the cosine similarity ($d_{bio}$ and $d_{app}$) and the Mahalanobis distance ($d_{pos}$). To avoid uncertain matches, $C_{i,j}^{app/bio}$ is gated by a maximum spatial distance $\theta_{pos}$, i.e.,
$C_{i,j}^{app/bio}=\begin{cases}\text{inf}&d_{pos}(\Phi_{i}^{bio},\Psi_{j}^{bio})>\theta_{pos}\\ C_{i,j}^{app/bio}&\text{otherwise}\end{cases},$ (3)
and a general threshold $\theta$ is applied, i.e.,
@@
Figure 1: Overview FaceQSORT association.
In DeepSORT, a matching cascade is applied in which the matching is performed in sequence, starting with the tracks that were correctly matched in the previous frame and ending with the tracks that have not been matched for the longest time. The matching cascade reduces the search space and could be beneficial when long occlusions are allowed. An overview of the core association procedure of FaceQSORT is depicted in Figure 1. For all unmatched detections (after the matching cascade), an IoU-based association (as proposed in SORT [15]) is performed as a fallback routine. The general steps that are performed in FaceQSORT for each frame are summarized in Algorithm 1.
@@
6:         calculate $C$ by Eqn. (1) using
7: $\Phi^{+age=i}$ and $\Psi^{\prime}$;
8:         find matches Eqn. (5)
9:         update $\Psi^{\prime}$ (unmatched detections)
10:     else
11:         break
12:     end if
13:end for
14:IoU fallback matching if $|\Psi^{\prime}|>0$
15:update/add $\Phi^{bio}$ and $\Phi^{app}$ for all detections (Eqn. (2))

### v3

@@ Abstract
In this work, a novel multi-face tracking method named FaceQSORT is proposed. To mitigate multi-face tracking challenges (e.g., partially occluded or lateral faces), FaceQSORT combines biometric and visual appearance features (extracted from the same image (face) patch) for association. The Q in FaceQSORT refers to the scenario for which FaceQSORT is desinged, i.e. tracking people’s faces as they move towards a gate in a Queue. This scenario is also reflected in the new dataset ‘Paris Lodron University Salzburg Faces in a Queue’, which is made publicly available as part of this work. The dataset consists of a total of seven fully annotated and challenging sequences (12730 frames) and is utilized together with two other publicly available datasets for the experimental evaluation. It is shown that FaceQSORT outperforms state-of-the-art trackers in the considered scenario. To provide a deeper insight into FaceQSORT, comprehensive experiments are conducted evaluating the parameter selection, a different similarity metric and the utilized face recognition model (used to extract biometric features).
@@ 1 Introduction
Multi-face tracking is about finding the trajectories of all faces (identities) appearing in a sequence of images. This work focuses on tracking the people’s faces while they move towards a gate in a queue.
@@
The assumed scenario is that a group of people is moving towards a gate (e.g., to enter a sports stadium). As they move towards a single gate, they form a queue. However, as this is a leisure activity, this queue will not be very well organized (people chat with each other, they eat, they move around in a disorderly fashion, there may be pushing and shoving, etc.). In the assumed scenario, the main objective is to track people’s faces as they move towards the gate. For this purpose, a camera is mounted on the gate. The camera is directed at the queue such that the first person visible is standing directly in front of the gate. However, since the queue is probably not very well organized, the main tracking challenges are occlusions (partially and full), out-of-plane rotations (lateral faces) and non-linear motion. In the assumed scenario, tracking people’s faces can facilitate entrance by enabling people who have registered a face image during pre-enrollment to pass through the gate without additional checks. On the other hand, it can help to identify unwanted persons (i.e., persons banned from the premises) on their way to the gate and deny them entrance. Furthermore, each ticket is usually personalized. Tracking people’s faces therefore enables a set of reference face images to be stored with the ticket, which assists in identifying individuals during the event.
@@
In this work, a novel multi-face tracker named FaceQSORT is proposed. FaceQSORT is designed to track multiple face in the assumed scenario (i.e., people are moving towards a gate in a queue). For this purpose, two different types of features are extracted from the same modality (i.e., the detected face). The two features (i.e., biometric and appearance features) are combined into a single cost value, which forms the basis for association and is the core of FaceQSORT. To the best of our knowledge, FaceQSORT is the first multi-face tracking method that combines two features extracted from the same image patch. In general, multi-face trackers [1, 2, 3, 4, 5, 6, 7] rely solely on biometric features. The combination of the two features helps to mitigate tracking challenges (e.g., lateral or partially occluded faces). FaceQSORT is evaluated based on a new multi-face tracking dataset. This dataset reflects the assumed scenario and is released as part of this work. To the best of our knowledge, the published dataset is the first multi-face tracking dataset in which people move in an (unorganized) queue towards a camera (gate). There is only one similar dataset, the ChokePoint dataset [8] (people moving through a portal), which is also used for experimental evaluation. In the released dataset, however, a queue is simulated in which people are talking, eating, pushing, shoving, etc. The third dataset used for experimental evaluation is the MusicVideo dataset [9]. This dataset consists of unconstrained videos that do not reflect the considered scenario and is only utilized when comparing FaceQSORT with state-of-the-art trackers.
@@
  • A novel multi-face tracking method (FaceQSORT) is proposed that combines two different features (i.e., biometric features and appearance features) extracted from the same image (face) patch.
  • A comprehensive experimental evaluation is performed on three datasets, comparing different face recognition models, analyzing parameter selection, applying a different similarity metric and comparing FaceQSORT with different state-of-the-art trackers.
@@
The remainder of this paper is organized as follows: related work is listed in section 2 and the proposed FaceQSORT is described in section 3. Section 4 provides an overview of the proposed dataset. The experimental setup is described in section 5. In section 6 the experimental results obtained are reported and discussed. The key insights are summarized in the last section 7.
2 Related Work
To solve a general MOT problem, basically two paradigms exist: (i) joint-detection-association (JDA), and (ii) tracking-by-detection (TbD). With a JDA based tracker, basically everything is learned (end-to-end), i.e. the object detection, the relevant cues for the association and the association of objects across frames. Recent methods in this context are [10, 11, 12, 13]. In this work, however, the focus is on TbD methods [14, 15, 16, 17]. The first step of a tracker based on the TbD paradigm is the detection of all objects (e.g., faces) in the current frame. The tracking performance is therefore directly related to the detection performance. Tracking is then about finding bipartite matches (i.e., association problem) for all detected objects and active tracks (i.e., detected objects from previous frames). A common way to solve this association problem is to apply the Hungarian algorithm [18] to a cost matrix (representing the costs of matching the $i$-th active track with the $j$-th detected object). These costs can, for example, be based on motion cues and/or appearance cues. Prominent examples for TbD trackers are Simple Online Realtime Tracking (SORT) [19] (motion based association), DeepSORT [20] (associated based on appearance features extracted by a Convolutional Neural Network (CNN)) and StrongSORT [21]. StrongSORT is an improved version of DeepSORT that comprises a more discriminative feature extractor [22], replaces the matching cascade with a vanilla matching, etc. A combination of motion and appearance models in a single network, called UMA, is proposed in [23]. In [24], a fusion of six distance metrics (for motion and appearance cues) is used for association. One advantage of TbD trackers over JDA trackers is that pre-trained models can be utilized for the individual subtasks (e.g., a face detector and a face recognition model), while performance is similar (as demonstrated in [25]).
In TbD based multi-face trackers, association is usually based on the similarity of biometric features (i.e., features extracted by a face recognition model). For example, a two-stage association based on motion and biometric features (ArcFace [26] trained on the MS-Celeb-1M dataset [27]) is proposed in [1]. Multi-face tracking based on DeepSORT [20] is proposed in [2]. The authors evaluate different combinations of face detectors (i.e., MTCNN [28], SSD [29] and R-FCN [30]) and a face recognition model trained with two different loss functions (i.e., cosine softmax classifier loss [31] and angle softmax classifier loss [32]). Biometric features of the upper facial regions (i.e., eyes, eyebrows and forehead) are used to track masked faces in [3]. A multi-face tracking method based on SORT [19] (i.e., motion based association) with a similarity matching block (comparing the stored with the currently computed biometric features, i.e. ArcFace [26]) as a fallback for all unmatched detections showing a frontal face is proposed in [4]. The method is named ReSORT, because IDs can be recovered based on the similarity matching block. Biometric features as fallback if motion based association fails is also proposed in [5]. In [6], a single object tracking method is used to predict the positions of an active track. The position-based matches are then corrected using biometric features. A multi-face tracking method based on biometric features only is presented in [7]. Cumulative learning of biometric features is performed based on a memory module, where selectively redundant features are removed. The proposed method is referred as IdOL (Identity Online Learning).
The biometric features used for tracking are usually based on pre-trained face recognition models (e.g., [26, 33, 34, 35]) that are typically trained on more or less frontal, unoccluded faces. In a multi-face tracking scenario, however, lateral faces (out-of-plane rotations) and partially occluded faces occur. To address this issue, a multi-modality tracker, Online Multi-face Tracking with Multi-modality Cascaded Matching (OMTMCM), is proposed in [36]. The proposed method utilizes body and face features and includes two stages: (i) detection alignment, and (ii) detection association. The detected faces and bodies are aligned in the first stage (i.e., detection alignment). Detection association is performed in a cascade, whereby matching based on biometric features is performed only with those detections that could not be matched with body features. The biometric features are extracted by a SeNet [37] pre-trained on the VGGFace2 database [38] and the body features by a ResNet [39] pre-trained on the MovieNet database [40]. However, a corresponding body is not always visible to every detected face, especially in the assumed scenario in which people are moving towards a gate in a queue.
Figure 1: The core of FaceQSORT. Biometric and appearance features are extracted from the same image (face) patch and the cosine similarity is calculated with all stored representations (features) from the active track memory $\Phi$.
@@ 3 FaceQSORT
FaceQSORT is a TbD-based tracker designed to track the Faces of multiple people as they move towards a gate in a Queue (source code is made available [^1] https://gitlab.cosy.sbg.ac.at/wavelab/faceqsort). SORT stands for Simple Online Realtime Tracking and illustrates that FaceQSORT is inspired by StrongSORT [21]. The core of FaceQSORT is the combination of two different types of features (i.e., biometric and appearance features) extracted from the same image (face) patch (i.e., the Bounding Box (BBox) returned by the face detector). By using two different features from the same modality, the required presence of body and face (as utilized in [36]), which hardly exists in the considered scenario, is eliminated. Furthermore, if only a single modality is utilized, the need for a second detector and some alignment overhead is avoided.
Biometric features, extracted by a face recognition model, provide a biometric representation of the face. These features should be distinct for each identity. However, due to tracking challenges (e.g., partially occluded, lateral or deformed faces), the distinctiveness of these features might be limited. For this reason, biometric features are combined with appearance features. Appearance features, extracted by a generic image classifier, provide a general visual representation of the image (face) patch and the combination with biometric features should mitigate multi-face tracking challenges. The combination of these features is performed on the cost matrix level and is depicted in Figure 1.
In common TbD based methods, the association relies on a cost matrix $C^{|\Phi|\times|\Psi|}$, where $\Phi$ is the set of active tracks (active track memory) and $\Psi$ the set of detected objects. An element $C_{i,j}$ represents the cost of matching the $i$-th active track to the $j$-th detected object. The main objective of association is to find bipartite matches where the global cost is minimum (i.e., usually solved by the Hungarian algorithm [18]). In principle, $C_{i,j}$ is based on a distance $d$ from a stored/predicted cue of the $i$-th active track to a calculated/observed cue of the $j$-th detected object. In FaceQSORT, the combination of biometric and appearance features is performed as follows,
@@
The distance measure $d_{bio}$ and $d_{app}$ is the cosine similarity (see Figure 1) of the stored and observed features and $\lambda\in\mathbb{R}_{[0,1]}$ weights (scales) the resulting costs (distances). The combination of biometric and appearance is unique (to the best of our knowledge) and represents the core of FaceQSORT.
Similar to StrongSORT a single cost matrix element is gated by a maximum spatial distance $\theta_{pos}$, i.e.
$C_{i,j}^{app/bio}=\begin{cases}\text{inf}&d_{pos}(\Phi_{i}^{pos},\Psi_{j}^{pos})>\theta_{pos}\\ C_{i,j}^{app/bio}&\text{otherwise}\end{cases},$ (2)
where $d_{pos}$ is the Mahalanobis distance. This is done to avoid uncertain matches. The spatial distance costs, $C_{i,j}^{pos}=d_{pos}(\Phi_{i}^{pos},\Psi_{i}^{pos})$, are also integrated into the final cost value, i.e.
$C_{i,j}=\beta C_{i,j}^{app/bio}+(1-\beta)C_{i,j}^{pos},$ (3)
and a general threshold $\theta$ is applied, i.e.
@@
Unlike StrongSORT, the matching cascade proposed with DeepSORT [20] is reactivated. In the matching cascade, matching is performed in sequence, starting with the tracks that were correctly matched in the previous frame and ending with the tracks that have not been matched for the longest time. The matching cascade reduces the search space and could be beneficial when long occlusions are allowed. For all unmatched detections (after the matching cascade), an IoU-based association (as proposed in SORT [19]) is performed as a fallback routine.
The active track memory $\Phi$, stores a unique ID, the biometric features $\Phi^{bio}$, the appearance features $\Phi^{app}$ and the position $\Phi^{pos}$ for each active track. A track becomes inactive (is deleted from $\Phi$) if it could not be matched with a detected face in $N_{max}$ consecutive frames. When a new track is added to $\Phi$, it is in a tentative state and is only confirmed if the track could be matched in the next $N_{init}$ frames. Otherwise, the tentative track is deleted from $\Phi$. Similar to StrongSORT, an Exponential Moving Average (EMA) [41] approach is used to update the stored biometric and appearance features from the $i$-th matched track, i.e.
$\Phi_{i}^{t}=\alpha\Phi_{i}^{t-1}+(1-\alpha)\Psi_{i}^{t},$ (6)
where $\Psi_{i}^{t}$ represents the extracted features from the current frame $t$ of the face associated with track $i$ and the parameter $\alpha$ is a momentum term. As stated in [21], this updating strategy leverages the information of inter-frame feature changes and can mitigate detection noises. $\Phi$ is maintained separately for biometric and appearance features (i.e., $\Phi^{bio}$ and $\Phi^{app}$). The prediction of the position of the $i$-th active track ($\Phi_{i}^{pos}$) in the next frame is based on the NSA Kalman filter [42]. The general steps that are performed in FaceQSORT for each frame are summarized in Algorithm 1.
@@
6:         calculate $C$ by Eqn. (3) using $\Phi^{+age=i}$ and $\Psi^{\prime}$
7:         find matches Eqn. (5)
8:         update $\Psi^{\prime}$ (unmatched detections)
9:     else
10:         break
11:     end if
12:end for
13:IoU fallback matching if $|\Psi^{\prime}|>0$
14:update/add $\Phi^{bio}$ and $\Phi^{app}$ for all detections (Eqn. (6))